OpenStudy (anonymous):

How many electrons in an atom can share the quantum numbers n=3

OpenStudy (btaylor):

there is 3s (2 electrons), 3p (6 electrons), and 3d (10 electrons) 2 + 6 + 10 = 18 electrons w/ n=3

OpenStudy (anonymous):

18

OpenStudy (anonymous):

2+6+10=18

OpenStudy (anonymous):

Correct @BTaylor & @suryachandra , does this help you understand why @Babeangel2737?